# On stochastic forms of functional isoperimetric inequalities

### v1

@@ 2.3. Functional mixed volumes
$\|u\|_{\Pi^{\circ}f}=n\int_{0}^{+\infty}V(\{f\geq t\}[n-1],[0,u]\})\,\mathrm{d}t=\int_{0}^{+\infty}|P_{u^{\perp}}\{f\geq t\}|_{n-1}\,\mathrm{d}t.$
Hence, when $f\in W^{1,1}(\mathbb{R}^{n})$, one obtains using co-areas’s formula that
@@
The latter makes readily verifiable that Zhang’s affine Sobolev inequality can be restated as
$|\Pi^{\circ}f|\leq|\Pi^{\circ}f^{*}|.$
@@ 3. Main results
Given an integrable $p$-concave function $f:\mathbb{R}^{n}\to\mathbb{R}_{+}$, with $p\in\mathbb{R}\cup\{+\infty\}$, and $\{(X_{k},Z_{k})\}_{k=1}^{N}$ be independent random vectors uniformly distributed w.r.t. $f$. As pointed out in [36], it is possible to consider more general forms of the stochastic model introduced in (1.8) by using the $M$-sum. Specifically, let $C\subset\mathbb{R}^{N}$ be a convex body contained in the positive orthant. We can define the random convex sets $E_{N,p},H_{N,p}\subset\mathbb{R}^{n}\times\mathbb{R}_{+}$ as
@@
when $p>0$. Note that the sets defined in (1.4) and (1.5) (and hence the stochastic model build upon them) can be recovered by setting $C=\mathop{\mathrm{conv}}\nolimits\{e_{1},\dots,e_{N}\}$.
The proofs of both Theorem 1.1 and 1.2 essentially rely on checking that the corresponding functionals satisfy the conditions of Theorem 2.3. In this regard, for readers convenience, we split the first part of Theorem 1.1 proof’s in two lemmas. We start by adapting [36, Proposition 5.1] to the context of level sets.
@@
Let $N\geq n+1$, $\{(x_{i},z_{i})\}_{i=1}^{N}\subset(\mathbb{R}^{n}\times\mathbb{R}_{+})^{N}$ and $\{\lambda_{i}\}_{i=1}^{N}\subset\mathbb{R}$. Let $C\subset\mathbb{R}^{N}$ be a convex body contained in the positive orthant,
@@
Let $N\geq n+1$ and $\{(x_{i},z_{i})\}_{i=1}^{N}\subset(\mathbb{R}^{n}\times\mathbb{R}_{+})^{N}$. Then, fixed $z_{1},\dots z_{N}\in\mathbb{R}_{+}$, the function $F_{u,y};\mathbb{R}^{N}\to\mathbb{R}_{+}$ given by
@@
For the reader convenience we will use the notation $\bar{s}=(s_{1},\dots,s_{N})$. Let $\bar{s},\bar{s}^{\prime}$ and $\lambda\in(0,1)$. As before, we only give a proof of the case $p\leq 0$;
@@
$\oplus_{C}\bigl{(}\{R_{\rho_{i}(p)}(y_{i}-s_{i}u)\}_{i=1}^{N}\bigr{)}=R_{u}\Bigl{(}\oplus_{C}\bigl{(}\{R_{\rho_{i}(s)}(y_{i}+s_{i}u)\}_{i=1}^{N}\bigr{)}\Bigr{)}$
@@
$\mathbb{E}$ $\Bigl{[}W_{i}\Bigl{(}\Phi^{N}_{(X_{k},Z_{k})}\Bigr{)}\Bigr{]}$
$=\left(\int f\right)^{-1}\int_{(\mathbb{R}^{n})^{N}}\int_{(\mathbb{R}_{+})^{N}}W_{i}\Bigl{(}\Phi^{N}_{(x_{k},z_{k})}\Bigr{)}\prod_{i=1}^{n}\chi_{{}_{\{f\geq z_{i}\}}}(x_{i})\,\mathrm{d}\bar{z}\,\mathrm{d}x$
$=\left(\int f\right)^{-1}\int_{(\mathbb{R}_{+})^{N}}\int_{(\mathbb{R}^{n})^{N}}\left(\int_{0}^{+\infty}W_{i}(\{\Phi^{N}_{(x_{k},z_{k})}\geq t\})\,\mathrm{d}t\right)\prod_{i=1}^{n}\chi_{{}_{\{f\geq z_{i}\}}}(x_{i})\,\mathrm{d}x\,\mathrm{d}\bar{z}$
$=\left(\int f\right)^{-1}\int_{(\mathbb{R}_{+})^{N}}\int_{(\mathbb{R}^{n})^{N}}F_{\bar{z}}(x_{1},\dots,x_{N})\prod_{i=1}^{n}g_{i}(x_{i})\,\mathrm{d}x\,\mathrm{d}\bar{z},$
@@
and $g_{i}(x_{i})=\chi_{{}_{\{f\geq z_{i}\}}}(x_{i})$ for all $i=1,\dots,N$. Hence, Lemma 3.2 implies that $F_{\bar{z}}$ is in the conditions of Theorem 2.3 for all $\bar{z}\in(\mathbb{R}_{+})^{N}$. Thus,
@@
$=\int_{(\mathbb{R}^{n})^{N}}F_{\bar{z}}(x_{1},\dots,x_{N})\chi_{{}_{\{f\geq z_{i}\}^{*}}}(x_{i})\,\mathrm{d}x,$
@@
$\mathbb{E}$ $\left[W_{i}\Bigl{(}\Phi^{N}_{(X_{k},Z_{k})}\Bigr{)}\right]$
$=\left(\int f\right)^{-1}\int_{(\mathbb{R}^{n})^{N}}\int_{(\mathbb{R}_{+})^{N}}F_{\bar{z}}(x_{1},\dots,x_{N})\prod_{i=1}^{n}\chi_{{}_{\{f\geq z_{i}\}}}(x_{i})\,\mathrm{d}\bar{z}\,\mathrm{d}x$
$\geq\left(\int f^{*}\right)^{-1}\int_{(\mathbb{R}^{n})^{N}}\int_{(\mathbb{R}_{+})^{N}}F_{\bar{z}}(x_{1},\dots,x_{N})\prod_{i=1}^{n}\chi_{{}_{\{f\geq z_{i}\}}}^{*}(x_{i})\,\mathrm{d}\bar{z}\,\mathrm{d}x$
$=\mathbb{E}\left[W_{i}\Bigl{(}\Phi^{N}_{(X^{*}_{k},Z^{*}_{k})}\Bigr{)}\right]$
@@
$\mathbb{E}$ $\bigl{[}\nu\bigl{(}\Pi^{\circ}\Phi^{N}_{(X_{k},Z_{k})}\bigr{)}\bigr{]}$
@@
$F_{\bar{z},y,u}(\bar{s})$ $=\int_{u^{\perp}}\int_{\mathbb{R}}\chi_{{}_{\{\|w+ru\|_{\Pi^{\circ}\Phi^{N}_{(y_{i}+s_{i}u_{i},z_{i})}}\leq 1\}}}\psi(w+ru)\,\mathrm{d}r\,\mathrm{d}w$
$=\int_{u^{\perp}}\int_{\mathbb{R}}\chi_{{}_{\{nV(\Phi^{N}_{(y_{i}+s_{i}u,z_{i})}[n-1],\chi_{{}_{[0,w+ru]}})\leq 1\}}}\psi(w+ru)\,\mathrm{d}r\,\mathrm{d}w$
@@
where $\varphi_{w}:u^{\perp}\times(\mathbb{R})^{N}\to\mathbb{R}_{+}$ is the function given by
@@
$V$ $(\Phi^{N}_{(y_{i}+((1-\lambda)s_{i}+\lambda s_{i}^{\prime})u,z_{i})}[n-1],\chi_{{}_{[0,w+((1-\lambda)r+\lambda r^{\prime})}})$
@@
Now, using Fubini’s
$\mathbb{E}$ $\bigl{[}\nu\bigl{(}\Pi^{\circ}\Phi^{N}_{(X_{k},Z_{k})}\bigr{)}\bigr{]}$
@@
$\mathbb{E}$ $\bigl{[}\nu\bigl{(}\Pi^{\circ}\Phi^{N}_{(X_{k},Z_{k})}\bigr{)}\bigr{]}$
@@
$\mathbb{E}\Bigl{[}\nu\bigl{(}\Pi^{\circ}\Phi^{N}_{(X_{k},Z_{k})}\bigr{)}\Bigr{]}\leq\mathbb{E}\Bigl{[}\nu\bigl{(}\Pi^{\circ}\Phi^{N}_{(X^{*}_{k},Z^{*}_{k})}\bigr{)}\Bigl{]},$
@@
We finish this section by showing how Theorems 1.1 and 1.2 recovers their geometrical counterparts. Let $K\subset\mathbb{R}^{n}$ be a convex body and $\{(X_{k},Z_{k})\}_{i=1}^{N}$ be independent random vectors uniformly distributed w.r.t. $\text{hyp}(\chi_{{}_{K}})=K\times[0,1]$. Since $\chi_{{}_{K}}$ is a $(+\infty)$-concave function, its stochastic model is given by
@@
Note that (3.6) recovers (1.2) when $C=\mathop{\mathrm{conv}}\nolimits\{e_{1},\dots,e_{N}\}$. Although not explicitly proved, this generalization of (1.2) was pointed out in [31, Remark 4.4].
@@
which was proved in [34, Theorem 1.2].
@@ 4. Recovering deterministic results
We would like to remark that epi-convergence has been widely use in the literature, specially in the study of valuations on convex functions (see e.g. [10, 11, 30, 12] and the references therein). We will use the following lemma originally proved in [11, Lemma 5]:
@@
We will use the stochastic models presented in (1.4) and (1.5) to see how Theorems 1.1 and 1.2 recover their deterministic counterparts. Let $N\geq n+1$, $f:\mathbb{R}^{n}\to\mathbb{R}_{+}$ be an integrable $p$-concave function and $\{(X_{k},Z_{k})\}_{k=1}^{N}$ be independent random vectors distributed w.r.t. f. Recall that the stochastic model of $f$ is build upon the random sets $E_{N,p},H_{N,p}\subset\mathbb{R}^{n}\times\mathbb{R}_{+}$ given by
@@
$\lim_{\varepsilon\to 0^{+}}\nu(\Pi^{\circ}f_{\varepsilon})=\nu(\Pi^{\circ}f).$
Therefore it is enough to prove our claims for $f_{\varepsilon}$.
Furthermore, let $\{(X_{k},Z_{k})\}_{k=1}^{N}$ be independent random vectors distributed w.r.t. $f_{\varepsilon}$. It is straightforward to check that the random sets $E_{N,p}$ and $H_{N,p}$ are convex bodies. Thus, using for instance Kolmogorov 010-1 law and Borel- Cantelli lemma, one can check that
$E_{N,p}\overset{\delta^{H}}{\to}epi(f_{\varepsilon}^{p})\quad\mathrm{if}\quad p<0,$
$E_{N,p}\overset{\delta^{H}}{\to}epi(-\log f_{\varepsilon})\quad\mathrm{if}\quad p=0\quad\mathrm{and}$
$H_{N,p}\overset{\delta^{H}}{\to}epi(f_{\varepsilon}^{p})\quad\mathrm{if}\quad p>0$
@@
We start by proving that almost sure (a. s. for short) convergence holds when considering the functional quermassintegral of an stochastic model. Although not explicitly described, we work in an underlying probability space.
@@
We will focus on the case of $p=0$ as the rest is analogous. Let $N\geq n+1$ and $\{(X_{k},Z_{k})\}_{i=1}^{N}$ be independent random vectors distributed w.r.t $f_{\varepsilon}$. We have that $E_{N,p}\subset\mathbb{R}^{n+1}$ is a convex body and that
@@
$W_{i}(f_{\varepsilon})=\mathbb{E}$ $\Bigr{[}\lim_{N\to+\infty}W_{i}\bigl{(}\Phi^{N}_{(X_{k},Z_{k})}\bigr{)}\Bigl{]}$
$=\lim_{N\to+\infty}\mathbb{E}\Bigl{[}W_{i}\bigl{(}\Phi^{N}_{(X_{k},Z_{k})}\bigr{)}\Bigr{]}$
@@
$W_{i}(f_{\varepsilon})$ $=\lim_{N\to+\infty}\mathbb{E}\Bigl{[}W_{i}\bigl{(}\Phi^{N}_{(X_{k},Z_{k})}\bigr{)}\Bigr{]}$
$\geq\lim_{N\to+\infty}\mathbb{E}\Bigl{[}W_{i}\bigl{(}\Phi^{N}_{(X^{*}_{k},Z^{*}_{k})}\bigr{)}\Bigr{]}$
@@
We now move on seeing how Theorem 1.2 recovers its determinist counterpart. First we prove the following.
@@
Let $N\geq n+1$ and $\{(X_{k},Z_{k})\}_{i=1}^{N}$ be independent random vectors distributed w.r.t $f_{\varepsilon}$. As before, we will focus on the case of $p=0$. Moreover, since polarity is continuous w.r.t. the Hausdorff distance, the claim is equivalent to check that
@@
$\lim_{N\to+\infty}||h_{\Pi f_{\varepsilon}}-h_{\Phi^{N}_{(X_{k},Z_{k})}}\|_{\infty}=0$
@@
$\nu(\Pi^{\circ}f_{\varepsilon})$ $=\mathbb{E}\Bigl{[}\nu\bigl{(}\lim_{N\to+\infty}\Pi^{\circ}\Phi^{N}_{(X_{k},Z_{k})}\bigr{)}\Bigr{]}$
$=\mathbb{E}\Bigl{[}\lim_{N\to+\infty}\nu\bigl{(}\Pi^{\circ}\Phi^{N}_{(X_{k},Z_{k})}\bigr{)}\Bigl{]}$
$=\lim_{N\to+\infty}\mathbb{E}\Bigl{[}\nu\bigl{(}\Pi^{\circ}\Phi^{N}_{(X_{k},Z_{k})}\bigr{)}\Bigl{]}.$
@@
$\nu(\Pi^{\circ}f_{\varepsilon})$ $=\lim_{N\to+\infty}\mathbb{E}\Bigl{[}\nu\bigl{(}\Pi^{\circ}\Phi^{N}_{(X_{k},Z_{k})}\bigr{)}\Bigl{]}$
$\leq\lim_{N\to+\infty}\mathbb{E}\Bigl{[}\nu\bigl{(}\Pi^{\circ}\Phi^{N}_{(X^{*}_{k},Z^{*}_{k})}\bigr{)}\Bigl{]}$
@@
Acknowledgments: We would like to thank Peter Pivovarov for helpful discussions and correspondence.

### v2

@@ 2.3. Functional mixed volumes
$\|u\|_{\Pi^{\circ}f}=n\int_{0}^{+\infty}V(\{f\geq t\}[n-1],[0,u])\,\mathrm{d}t=\int_{0}^{+\infty}|P_{u^{\perp}}\{f\geq t\}|_{n-1}\,\mathrm{d}t.$
Hence, when $f$ belongs to the Sobolev space $W^{1,1}(\mathbb{R}^{n})$ of integrable functions with integrable first-order weak derivatives, one obtains using the co-area formula that
@@
From the latter, it is not difficult to check that the inequality
$|\Pi^{\circ}f|\leq|\Pi^{\circ}f^{*}|,$
is equivalent to Zhang’s affine Sobolev inequality restricted to the class of $p$-concave functions in $W^{1,1}(\mathbb{R}^{n})$.
@@ 3. Main results
Given an integrable $p$-concave function $f:\mathbb{R}^{n}\to\mathbb{R}_{+}$, with $p\in\mathbb{R}\cup\{+\infty\}$, and let $\{(X_{k},Z_{k})\}_{k=1}^{N}$ be independent random vectors uniformly distributed w.r.t. $f$. As pointed out in [36], it is possible to consider more general forms of the stochastic model introduced in (1.11) by using the $M$-sum. Specifically, let $C\subset\mathbb{R}^{N}$ be a compact convex set contained in the positive orthant. We can define the random convex sets $E_{N,p},H_{N,p}\subset\mathbb{R}^{n}\times\mathbb{R}_{+}$ as
@@
when $p>0$. Note that the sets defined in (1.4) and (1.5) (and hence the stochastic model built upon them) can be recovered by setting $C=\mathop{\mathrm{conv}}\nolimits\{e_{1},\dots,e_{N}\}$.
The proofs of both Theorem 1.1 and 1.2 essentially rely on checking that the corresponding functionals satisfy the conditions of Theorem 2.3. In this regard, for the reader’s convenience, we split the first part of Theorem 1.1 proof’s in two lemmas. We start by adapting [36, Proposition 5.1] to the context of level sets.
@@
Let $N\geq n+1$, $\{(x_{i},z_{i})\}_{i=1}^{N}\subset(\mathbb{R}^{n}\times\mathbb{R}_{+})^{N}$ and $\{\lambda_{i}\}_{i=1}^{N}\subset\mathbb{R}$. Let $C\subset\mathbb{R}^{N}$ be a compact convex set contained in the positive orthant,
@@
Let $N\geq n+1$ and $\{(x_{i},z_{i})\}_{i=1}^{N}\subset(\mathbb{R}^{n}\times\mathbb{R}_{+})^{N}$. Then, for fixed $z_{1},\dots z_{N}\in\mathbb{R}_{+}$, the function $F_{u,y};\mathbb{R}^{N}\to\mathbb{R}_{+}$ given by
@@
For the reader’s convenience we will use the notation $\bar{s}=(s_{1},\dots,s_{N})$. Let $\bar{s},\bar{s}^{\prime}$ and $\lambda\in(0,1)$. As before, we only give a proof of the case $p\leq 0$; note that
@@
$\oplus_{C}\bigl(\{R_{\rho_{i}(p)}(y_{i}-s_{i}u)\}_{i=1}^{N}\bigr)=R_{u}\Bigl(\oplus_{C}\bigl(\{R_{\rho_{i}(p)}(y_{i}+s_{i}u)\}_{i=1}^{N}\bigr)\Bigr)$
@@
[…]
$=\left(\int f\right)^{-N}\int_{(\mathbb{R}^{n})^{N}}\int_{(\mathbb{R}_{+})^{N}}W_{i}\Bigl(\Phi^{N}_{(x_{k},z_{k})}\Bigr)\prod_{i=1}^{N}\chi_{{}_{\{f\geq z_{i}\}}}(x_{i})\,\mathrm{d}\bar{z}\,\mathrm{d}x$
$=\left(\int f\right)^{-N}\int_{(\mathbb{R}_{+})^{N}}\int_{(\mathbb{R}^{n})^{N}}\left(\int_{0}^{+\infty}W_{i}(\{\Phi^{N}_{(x_{k},z_{k})}\geq t\})\,\mathrm{d}t\right)\prod_{i=1}^{N}\chi_{{}_{\{f\geq z_{i}\}}}(x_{i})\,\mathrm{d}x\,\mathrm{d}\bar{z}$
@@
and $g_{i}(x_{i})=\chi_{{}_{\{f\geq z_{i}\}}}(x_{i})$ for all $i=1,\dots,N$. Hence, Lemma 3.2 implies that $F_{\bar{z}}$ satisfies the conditions of Theorem 2.3 (those of the quasi-convex analogue) for all $\bar{z}\in(\mathbb{R}_{+})^{N}$. Thus,
@@
$=\int_{(\mathbb{R}^{n})^{N}}F_{\bar{z}}(x_{1},\dots,x_{N})\prod_{i=1}^{N}\chi_{{}_{\{f\geq z_{i}\}^{*}}}(x_{i})\,\mathrm{d}x,$
@@
[…]
$=\left(\int f\right)^{-N}\int_{(\mathbb{R}^{n})^{N}}\int_{(\mathbb{R}_{+})^{N}}F_{\bar{z}}(x_{1},\dots,x_{N})\prod_{i=1}^{N}\chi_{{}_{\{f\geq z_{i}\}}}(x_{i})\,\mathrm{d}\bar{z}\,\mathrm{d}x$
$\geq\left(\int f^{*}\right)^{-N}\int_{(\mathbb{R}^{n})^{N}}\int_{(\mathbb{R}_{+})^{N}}F_{\bar{z}}(x_{1},\dots,x_{N})\prod_{i=1}^{N}\chi_{{}_{\{f\geq z_{i}\}}}^{*}(x_{i})\,\mathrm{d}\bar{z}\,\mathrm{d}x$
[…]
@@
[…]
@@
[…]
$=\int_{u^{\perp}}\int_{\mathbb{R}}\bigl[w+ru\in\{\|w+ru\|_{\Pi^{\circ}\Phi^{N}(y_{i}+s_{i}u_{i},z_{i})}\leq 1\}\bigr]\psi(w+ru)\,\mathrm{d}r\,\mathrm{d}w$
$=\int_{u^{\perp}}\int_{\mathbb{R}}\bigl[w+ru\in\{nV(\Phi^{N}(y_{i}+s_{i}u,z_{i})[n-1],\chi_{{}_{[0,w+ru]}})\leq 1\}\bigr]\psi(w+ru)\,\mathrm{d}r\,\mathrm{d}w$
@@
where, when needed, we use $x\in A]$ for $\chi_{{}_{A}}$ to make reading easier. Moreover, $\varphi_{w}:u^{\perp}\times(\mathbb{R})^{N}\to\mathbb{R}_{+}$ is the function given by
@@
$V$ $(\Phi^{N}_{(y_{i}+((1-\lambda)s_{i}+\lambda s_{i}^{\prime})u,z_{i})}[n-1],\chi_{{}_{[0,w+((1-\lambda)r+\lambda r^{\prime})u]}})$
@@
Now, using Fubini’s theorem
[…]
@@
[…]
@@
[…]
@@
We finish this section by showing how Theorems 1.1 and 1.2 recover their geometrical counterparts. Let $K\subset\mathbb{R}^{n}$ be a convex body and $\{(X_{k},Z_{k})\}_{k=1}^{N}$ be independent random vectors uniformly distributed w.r.t. $\text{hyp}(\chi_{{}_{K}})=K\times[0,1]$. Since $\chi_{{}_{K}}$ is a $(+\infty)$-concave function, its stochastic model is given by
@@
Note that (3.6) recovers (1.2) when $C=\mathop{\mathrm{conv}}\nolimits\{e_{1},\dots,e_{N}\}$. Although not explicitly proved, this generalization of (1.2) was pointed out in [30, Remark 4.4].
@@
which was proved in [33, Theorem 1.2].
@@ 4. Recovering deterministic results
We would like to remark that epi-convergence has been widely used in the literature, especially in the study of valuations on convex functions (see e.g. [10, 11, 29, 12] and the references therein). We will use the following lemma originally proved in [11, Lemma 5]:
@@
We will use the stochastic models presented in (1.4) and (1.5) to see how Theorems 1.1 and 1.2 recover their deterministic counterparts. Let $N\geq n+1$, $f:\mathbb{R}^{n}\to\mathbb{R}_{+}$ be an integrable $p$-concave function and $\{(X_{k},Z_{k})\}_{k=1}^{N}$ be independent random vectors distributed w.r.t. f. Recall that the stochastic model of $f$ is built upon the random sets $E_{N,p},H_{N,p}\subset\mathbb{R}^{n}\times\mathbb{R}_{+}$ given by
@@
$\lim_{\varepsilon\to 0^{+}}\nu(\Pi^{\circ}f_{\varepsilon})=\nu(\Pi^{\circ}f),$
where we have used that the functional polar projection body is continuous (see [10, Lemma 4.1]). Therefore it is enough to prove our claims for $f_{\varepsilon}$.
Furthermore, let $\{(X_{k},Z_{k})\}_{k=1}^{N}$ be independent random vectors distributed w.r.t. $f_{\varepsilon}$. It is straightforward to check that the random sets $E_{N,p}$ and $H_{N,p}$ are convex bodies. Thus, using for instance Kolmogorov’s 010-1 law and Borel- Cantelli lemma, one can check that
$E_{N,p}\overset{\delta^{H}}{\to}\mathrm{epi}(f_{\varepsilon}^{p})\quad\mathrm{if}\quad p<0,$
$E_{N,p}\overset{\delta^{H}}{\to}\mathrm{epi}(-\log f_{\varepsilon})\quad\mathrm{if}\quad p=0\quad\mathrm{and}$
$H_{N,p}\overset{\delta^{H}}{\to}\mathrm{hyp}(f_{\varepsilon}^{p})\quad\mathrm{if}\quad p>0$
@@
We start by proving that almost sure (a. s. for short) convergence holds when considering the functional quermassintegral of a stochastic model. Although not explicitly described, we work in an underlying probability space.
@@
We will focus on the case of $p=0$ as the rest is analogous. Let $N\geq n+1$ and $\{(X_{k},Z_{k})\}_{k=1}^{N}$ be independent random vectors distributed w.r.t $f_{\varepsilon}$. We have that $E_{N,p}\subset\mathbb{R}^{n+1}$ is a convex body and that
@@
[…]
@@
[…]
@@
We now move on seeing how Theorem 1.2 recovers its deterministic counterpart. First we prove the following.
@@
Let $N\geq n+1$ and $\{(X_{k},Z_{k})\}_{i=1}^{N}$ be independent random vectors distributed w.r.t $f_{\varepsilon}$. As before, we will focus on the case of $p=0$. Moreover, since polarity is continuous w.r.t. the Hausdorff distance for convergence of centrally symmetric convex bodies, the claim is equivalent to check that
@@
$\lim_{N\to+\infty}\|h_{\Pi f_{\varepsilon}}-h_{\Phi^{N}_{(X_{k},Z_{k})}}\|_{\infty}=0$
@@
[…]
@@
[…]
@@
Acknowledgments: We would like to thank Peter Pivovarov for helpful correspondence, and Fabian Mussnig for valuable comments. We also extend our sincere gratitude to the anonymous referee, whose remarks have greatly enhanced the clarity and quality of our manuscript.
Conflicts of Interest: There are no conflicts of interest.